# WavLLM: Towards Robust and Adaptive Speech Large Language Model

### v2

@@ Abstract
The recent advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), progressively broadening their scope to multimodal perception and generation. However, effectively integrating listening capabilities into LLMs poses significant challenges, particularly with respect to generalizing across varied contexts and executing complex auditory tasks. In this work, we introduce WavLLM, a robust and adaptive speech large language model with dual encoders, and a prompt-aware LoRA weight adapter, optimized by a two-stage curriculum learning approach. Leveraging dual encoders, we decouple different types of speech information, utilizing a Whisper encoder to process the semantic content of speech, and a WavLM encoder to capture the unique characteristics of the speaker’s identity. Within the curriculum learning framework, WavLLM first builds its foundational capabilities by optimizing on mixed elementary single tasks, including automatic speech recognition (ASR), speech translation (ST), speaker verification (SV), emotion recognition (ER), instruction tuning (IT) and speech question answering (SQA), followed by advanced multi-task training on more complex tasks such as combinations of the elementary tasks. To enhance the flexibility and adherence to different tasks and instructions, a prompt-aware LoRA weight adapter is introduced in the second advanced multi-task training stage. We validate the proposed model on universal speech benchmarks including tasks such as ASR, ST, SV, ER, and also apply it to specialized datasets like Gaokao English listening comprehension set for SQA, and speech Chain-of-Thought (CoT) evaluation set. Experiments demonstrate that the proposed model achieves state-of-the-art performance across a range of speech tasks on the same model size, exhibiting robust generalization capabilities in executing complex tasks using CoT approach. Furthermore, our model successfully completes Gaokao English listening comprehension tasks without specialized training. The codes, models, audio samples, and Gaokao evaluation set can be accessed at aka.ms/wavllm.
@@ 1 Introduction
Large language models (LLMs) have witnessed a meteoric rise in advancement within the last couple of years, reaching or even exceeding the proficiency of humans in a myriad of natural language processing (NLP) tasks (OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023). These impressive capabilities are attributable to the large-scale training of vast training datasets and a substantial number of model parameters, combined with advanced training methodologies, like instruction-following protocols and Reinforcement Learning from Human Feedback (RLHF) algorithms (Ouyang et al., 2022). With large language models attaining substantial breakthroughs, the focus is increasingly shifting towards the capabilities and advancements of multi-modal large language models (MLLMs), which possess the ability to listen (Tang et al., 2023; Deshmukh et al., 2023), speak (Rubenstein et al., 2023; Hao et al., 2023), see (Huang et al., 2023; OpenAI, 2023), and create content (Pan et al., 2023; Brooks et al., 2024).
@@
Amidst the broadening scope of abilities, speech stands out as a crucial form of human communication, prompting extensive research to equip large language models (LLMs) with speech perception capabilities (Shu et al., 2023; Wu et al., 2023; Tang et al., 2023; Chu et al., 2023; Ma et al., 2024). Typically, LLMs are augmented with an auxiliary audio encoder designed to preprocess audio signals, transforming them into the same input space as that of the LLMs, enabling them to achieve various speech tasks, such as automatic speech recognition (ASR), speech question answering (SQA), and so on. However, previous research has yet to overcome significant challenges in achieving effective generalization due to two main issues: (1) specialized tasks are highly sensitive to prompt design, resulting in performance degradation when confronted with unseen or complex instructions; (2) there is an absence of speech Chain-of-Thought (CoT) capability, which is essential for addressing complex tasks.
@@
In this work, we propose a robust and adaptive speech large language model, WavLLM, aiming at enhancing the generalization capabilies, following speech instruction effectively, and processing the given speech in accordance with provided textual prompts, as well as supporting multi-round dialog. Specifically, to distinguish various types of speech information, we utilize a Whisper (Radford et al., 2023) encoder to encode the semantic content of the speech, and a WavLM (Chen et al., 2022) encoder to capture the acoustic information, like unique characteristics of the speaker’s identity. During the model training phase, we develop a curriculum learning method that progressively fine-tune LLMs to follow instructions for understanding and processing speech, starting from simple tasks and advancing towards more complex ones. In the initial mixed single-task training stage, we leverage a substantial dataset of synthesized spoken question-answering content generated by GPT-4 and tailored to various speech-centric tasks such as automatic speech recognition (ASR), speech-to-text translation (ST), emotion recognition (ER), speaker verification (SV), and so on, to fine-tune the WavLLM with Low Rank Adaptation (LoRA) techniques (Hu et al., 2021). Our preliminary experiments indicate that incorporating speech modality into LLMs by efficient fine-tuning technique, e.g., LoRA, may result in over-fitting on the training tasks, particularly on ASR task, and leading to subpar generalization on the unseen or complex instructions. To this end, we introduce a novel prompt-aware LoRA weight adapter in the advanced multi-task training stage. This adapter, trained on a specially constructed prompt-aware multi-task dataset, such as combinations of the elementary tasks, can adaptively adjust the LoRA weights in response to the different prompts, with the goal of enhancing the generalization ability of the initial model.
@@
We evaluate the proposed model on 1) single tasks, including a) universal speech benchmark, for speech recognition, speaker verification, emotion recognition and speech translation; b) spoken-query-based question answering and English Listening Comprehension test in Chinese National College Entrance Examination (Gaokao), which presents a spoken dialogue between various speakers, and requires to answer text-based single-answer multiple-choice questions related to the conversation, and 2) multiple tasks, consisting of c) instruction-independent multi-tasks dataset that combines multiple independent prompts within a single instruction; d) speech CoT evaluation set that decomposes a complex task into multiple sub-tasks for effective assessment. Extensive evaluations demonstrate that our proposed model exhibits robust generalization and CoT capabilities, consistently surpassing strong baselines across a broad spectrum of speech-related tasks.
@@
  • Equipped with a prompt-aware LoRA weight adapter, we introduce a curriculum learning-based training approach that progressively fine-tunes large language models to follow instructions with strong speech processing and comprehension capabilities, starting from simple tasks and advancing to more complex ones.
  • Our model decouples speech information into semantic and acoustic components, employing the Whisper encoder and the WavLM encoder to model them separately, which enhances the speech representation and yields benefits for downstream tasks.
  • WavLLM demonstrates exceptional generalization capabilities when responding to diverse prompts and completing complex tasks. It exhibits impressive capabilities in zero-shot SQA such as Gaokao English listening comprehension, and shows strong proficiency in CoT-based tasks, delivering performance gains over non-CoT tasks.
@@ 2 Related Work
The exploration of multi-modal large language models involves the integration of diverse data types including text, images, video, speech, audio, and more. This represents a natural progression from text-based large language models, designed to enable the perception of the world and the creation of content through the integration of various forms of input beyond text (OpenAI, 2023; Huang et al., 2023; Hao et al., 2023). For instance, Kosmos-1 (Huang et al., 2023) and GPT-4V (OpenAI, 2023) are able to perceive general modalities beyond text, and follow instruction provided by users to process and analyze image inputs. Another research direction focuses on improving the multi-modal generative abilities of language models, enabling them to produce visual content like images or videos, as exemplified by MiniGPT-5 (Zheng et al., 2023) and Sora (Brooks et al., 2024). Related research to this work focuses on speech-enhanced large language models that aim to endow LLMs with the capability to perceive and process speech signal (Zhang et al., 2023; Shu et al., 2023; Wu et al., 2023; Tang et al., 2023; Chu et al., 2023).
Among these studies, SpeechGPT (Zhang et al., 2023) empowers large language models with cross-modal conversational abilities by three-stage training stages, using hidden units as the discrete representation of speech. It constructs a speech-text cross-modal instruction-following dataset based on ASR data and textual SFT data. LLaSM (Shu et al., 2023) builds a large Chinese/English speech-language model that can understand and follow instructions, through pre-training stage and cross-modal instruction fine-tuning stage. SALMONN (Tang et al., 2023), named from a speech audio language music open neural network, boosts large language models with generic hearing abilities with a activation tuning stage by playing with the LoRA scaling factor. Qwen-audio (Chu et al., 2023) scales up audio-language pre-training to cover over 30 tasks and various audio types, including human speech, natural sounds, music, and songs. Despite achieving impressive results on specific speech tasks and being capable of following instructions to complete tasks, they continue to encounter challenges with regard to their weak generalization capabilities to complex tasks.
@@
Previous research on Speech Large Language Models (Speech LLMs) has primarily concentrated on executing a single speech task in response to a given instruction, while the feasibility of using a single instruction to simultaneously complete multiple and complex speech tasks has remained unexplored. The employment of multi-task instructions allows for the efficient completion of several tasks at once and improves performance by dividing complex tasks into logical, related sub-tasks, such as CoT tasks.
Our initial experiments indicate that (1) prior open-source speech LLMs underperformed in multi-task scenarios, demonstrating these models’ limited ability to generalize to complex instructions; (2) reducing the LoRA scaling factor can be beneficial for multi-task instructions, but leads to a substantial degradation of the results of training tasks (Tang et al., 2023), which suggests that single and multiple tasks might benefit from distinct LoRA scaling factors; (3) there is a notable decline in performance when the model encounters unseen or diverse prompts as opposed to seen prompts (3.5% vs. 2.1%, see Section 4.4), when employing various prompts to evaluate the ASR performance of the open-source model. Consequently, we introduce a curriculum learning approach that progresses from simple to complex instruction tasks, propose a prompt-aware LoRA weight adapter which dynamically adjusts the amplitude of the LoRA output according to the instruction, and further enhance the generalization by utilizing a diverse array of prompts generated by GPT-4 across all training tasks.
@@ 3 Method
In this section, we will first introduce the model architecture of the proposed framework (See subsection 3.1), consisting of dual speech encoders with modality adapters, and LLM with prompt-aware LoRA weight adapter. In subsection 3.2, we outline the training stages based on curriculum learning, which initially equip the LLM with strong speech processing and comprehension capabilities. Subsequently, in the second advanced multi-task training stage, the model is augmented with the capacity to generalize across complex multi-task instructions, achieving this without compromising its performance on single speech tasks. The WavLLM is optimized by maximizing the following probability:
@@
where $\bm{X}$ and $\bm{T}$ are the speech input and text prompt respectively. $\bm{Y}=[\bm{y}_{1},\bm{y}_{2},...,\bm{y}_{T_{Y}}]$ is the target text output. $\Theta$ denotes the parameters of WavLLM. The detailed template of WavLLM’s training data can be found in Appendix D.
@@ 3.1 Model Architecture
The model architecture of our framework is shown in Figure 1, which consists of speech encoders (i.e., Whisper (Radford et al., 2023) and WavLM (Chen et al., 2022)) as well as modality adapters, a large language model (i.e., LLaMA (Touvron et al., 2023)) and a proposed prompt adapter.
@@
In order to extract both the semantic and acoustic information in the speech, we utilize two state-of-the-art speech encoders, namely Whisper and WavLM. Whisper is trained for ASR and ST tasks in a weakly supervised fashion on a massive 680k-hour speech corpus recorded in diverse conditions, making it well suited for encoding semantic information in speech. WavLM is a predictive based self-supervised learning (SSL) pre-trained model. During its pre-training stage, WavLM mixes each utterance with signals from multiple speakers in the same batch, yet selectively predicts only the targets associated with the utterance’s original speaker. Such training method allows WavLM to better extract speaker-related acoustic information. In our work, the 32-layer transformer-based encoder of Whisper-large-v2 and WavLM-base are utilized. Both modality adapters have three components, including two 1-D convolution layers to down-sample and align the output of both encoders within the temporal domain, a down-up bottleneck adapter (Houlsby et al., 2019), and a final linear projector. The semantic adapter receives its input from the Whisper encoder’s output, while the acoustic adapter takes a weighted sum of the hidden states from all layers of WavLM, where the weights are learnable. The outputs of both adapters are concatenated together before feedforwarding into the linear projector.
@@
To enable adaptive LoRA scaling factors for different single-task and multiple-task instructions, inspired by adapter layer in Houlsby et al. (2019), we design an online adaptation strategy by introducing a down-up prompt-aware LoRA weight adapter (aka. prompt adapter) with attention weights, designed to modulate the effect of LoRA on LLaMA, as shown in Figure 1. This is achieved by producing $D$-dim tailored parameters in response to the provided instruction, prior to LoRA’s interaction with LLaMA, where $D$ is the dimension of the LoRA representations. As shown in Figure 1, given the text-based prompts $\bm{T}$ with length $M$, we can get the representation $\bm{t}\in\mathbb{R}^{D\times M}$ with LLAMA, $\bm{t}=f(\bm{T};\bm{\Theta_{\text{LLaMA}}})$, and this representation is fed into the prompt adapter to get the LoRA scaling factors, $\bm{r}\in\mathbb{R}^{D\times 1}$:
@@
where $\bm{W}_{A}\in\mathbb{R}^{1\times D}$ is the matrix of attention weights, $\bm{I}_{D}(\bm{W}_{A}\bm{o})\in\mathbb{R}^{D\times M}$ is unnormalized weights for the hidden states $\bm{o}\in\mathbb{R}^{D\times M}$. $\bm{I}_{D}\in\mathbb{R}^{D\times 1}$ and $\bm{I}_{M}\in\mathbb{R}^{M\times 1}$ are the all-ones vectors. $\bm{P}^{u}\in\mathbb{R}^{D\times K}$ and $\bm{P}^{d}\in\mathbb{R}^{K\times D}$ are up-linear projection and down-linear projection layers respectively, and GeLU is the GeLU activation function. The hidden states of an attention layer equipped with adaptive LoRA are expressed by:
@@ 3.2.1 Mixed Single-Task Training Stage
In the first stage, various single-task, cross-modal, speech-text pairing datasets or text-only datasets are utilized to endow the LLM with robust capabilities in speech processing and comprehension. We freeze the parameters of LLM, WavLM, and Whisper encoder, and optimize the modality adapters, a linear layer and LoRA components within the large language model.
@@
The first mixed single-task training stage involves various speech tasks, including automatic speech recognition (ASR), speech-to-text translation (ST), speaker verification (SV), emotion recognition (ER), spoken-based instruction tuning and text-based instruction tuning (IT) tasks, as well as a large mount of GPT-generated speech question answering (SQA). There are various questions within the SQA tasks, including those related to the speaker and gender, as well as continuation and summary tasks. Concurrently, these tasks draw upon multiple datasets, including LibriSpeech (Panayotov et al., 2015) with English reading speech, AMI (Carletta et al., 2005) with multi-talker meeting recordings, as well as Fisher (Cieri et al., 2004) and Switchboard (Godfrey et al., 1992) corpora with 2-channel English telephony conversations. Examples of the training data and prompts used to generate data with GPT-4 can be found in the Appendix A.1 and A.3 respectively. The speech audio clips of spoken-based instruction tuning task are generated by using Microsoft Azure text-to-speech API [^1] https://azure.microsoft.com/en-us/products/ai-services/text-to-speech. In Table 1, we list the detailed task information about description, data source, and data hours.

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have expanded their scope in natural language processing (NLP) to encompass multimodal functions. However, integrating listening capabilities effectively remains a significant challenge for generalization and complex auditory task execution. In this work, we introduce WavLLM, a robust and adaptive speech large language model featuring dual encoders—a Whisper encoder for semantics and a WavLM encoder for speaker characteristics. Within the two-stage curriculum learning framework, WavLLM first builds its foundational capabilities by optimizing on mixed elementary single tasks, followed by advanced multi-task training on more complex tasks combining elementary ones. To enhance the flexibility and adherence to different tasks and instructions, a prompt-aware LoRA weight adapter is introduced in the second advanced multi-task training stage. We validate the proposed model on universal speech benchmarks and also apply it to specialized speech-question-answer (SQA) dataset, and speech Chain-of-Thought (CoT) evaluation set. Experiments demonstrate that the proposed model achieves state-of-the-art performance across a range of speech tasks in the setting of the same model size, exhibiting robust generalization capabilities in executing complex tasks using CoT approach. The codes, models, audio samples, and SQA evaluation set can be accessed at https://github.com/wavllm/wavllm-anonymous.
WavLLM: Towards Robust and Adaptive Speech Large Language Model
Anonymous ACL submission
@@ 1 Introduction
Large language models (LLMs) have witnessed a meteoric rise in advancement within the last couple of years, reaching or even exceeding the proficiency of humans in a myriad of natural language processing (NLP) tasks (OpenAI2023GPT4TR; touvron2023llama; anil2023palm). With large language models attaining substantial breakthroughs, the focus is increasingly shifting towards the capabilities and advancements of multi-modal large language models (MLLMs), which possess the ability to listen (tang2023salmonn; deshmukh2023pengi), speak (rubenstein2023audiopalm; hao2023boosting), see (huang2023language; OpenAI2023GPT4TR), and create content (pan2023generating; videoworldsimulators2024).
@@
Amidst the broadening scope of abilities, speech stands out as a crucial form of human communication, prompting extensive research to equip large language models (LLMs) with speech perception capabilities (shu2023llasm; wu2023decoder; wang2023blsp; tang2023salmonn; chu2023qwen). Typically, LLMs are augmented with an auxiliary audio encoder designed to preprocess audio signals, transforming them into the same input space as that of the LLMs, enabling them to achieve various speech tasks, such as automatic speech recognition (ASR), speech question answering (SQA), and so on. However, previous research has yet to overcome significant challenges in achieving effective generalization due to two main issues: (1) specialized tasks are highly sensitive to prompt design, resulting in performance degradation when confronted with unseen or complex instructions; (2) there is an absence of speech Chain-of-Thought (CoT) wei2022chain capability, which is essential for addressing complex tasks.
In this work, we propose a robust and adaptive speech large language model, WavLLM, aiming at enhancing the generalization capabilies, following speech instruction effectively, and processing the given speech in accordance with provided textual prompts, as well as supporting multi-round dialog. Specifically, to distinguish various types of speech information, we utilize a Whisper (radford2023robust) encoder to encode the semantic content of the speech, and a WavLM (chen2022wavlm) encoder to capture the acoustic information, like unique characteristics of the speaker’s identity.
@@
During the model training phase, we develop a curriculum learning method that progressively fine-tune LLMs to follow instructions for understanding and processing speech, starting from simple tasks and advancing towards more complex ones. In the initial mixed single-task training stage, we leverage a substantial dataset of synthesized spoken question-answering content generated by GPT-4 and tailored to various speech-centric tasks such as automatic speech recognition (ASR), speech-to-text translation (ST), emotion recognition (ER), speaker verification (SV), and so on, to fine-tune the WavLLM with Low Rank Adaptation (LoRA) techniques (hu2022lora). To enhance the generalization on the unseen or complex instructions [^1] Please find detailed motivations in Section 2., we introduce an advanced multi-task training stage, incorporating a specially constructed prompt-aware multi-task speech processing dataset combining the elementary tasks. Furthermore, we design a novel prompt-aware LoRA weight adapter for this stage, capable of adaptively tuning the LoRA weights according to varying prompts, thereby improving the model’s generalization capacity and robustness.
@@
We evaluate the proposed model on 1) single tasks, including a) universal speech benchmark, including ASR, SV, ER and ST; b) spoken-query-based question answering and English Listening Comprehension test in Chinese National College Entrance Examination, which presents a spoken dialogue, and requires to answer text-based choice questions related to the conversation, and 2) multiple tasks, consisting of c) instruction-independent multi-tasks dataset that combines multiple independent prompts in a single instruction; d) speech CoT evaluation set that decomposes a complex task into multiple sub-tasks. Extensive evaluations demonstrate that our proposed model exhibits robust generalization and CoT capabilities, consistently surpassing strong baselines across a broad spectrum of speech-related tasks.
@@
1) Equipped with the proposed prompt-aware LoRA weight adapter, a curriculum learning method is leveraged for model training, by incrementally fine-tuning large language models with robustness and generalization capabilities, beginning with simple tasks and progressing to complex ones.
2) Our proposed model employs a decoupling strategy for speech information, utilizing the Whisper encoder to capture semantic content and the WavLM encoder for acoustic features, thereby enriching speech representation and improving performance on downstream tasks.
3)WavLLM demonstrates exceptional generalization capabilities when responding to diverse prompts and completing complex tasks. It exhibits impressive capabilities in zero-shot SQA such as English listening comprehension, and shows strong proficiency in CoT-based tasks, delivering performance gains over non-CoT tasks.
@@ 2 Related Work
The exploration of multi-modal large language models involves the integration of diverse data types including text, images, video, speech, audio, and more. This represents a natural progression from text-based large language models, designed to enable the perception of the world and the creation of content (OpenAI2023GPT4TR; huang2023language; hao2023boosting). For instance, Kosmos-1 (huang2023language) and GPT-4V (OpenAI2023GPT4TR) are able to perceive general modalities beyond text, and follow instruction provided by users to process and analyze image inputs. Another research direction focuses on improving the multi-modal generative abilities of language models, enabling them to produce visual content like images or videos, as exemplified by MiniGPT-5 (zheng2023minigpt) and Sora (videoworldsimulators2024). Related research to this work focuses on speech-enhanced large language models that aim to endow LLMs with the capability to perceive and process speech signal (zhang2023speechgpt; shu2023llasm; wu2023decoder; tang2023salmonn; chu2023qwen; wang2023blsp).
Among these studies, SpeechGPT (zhang2023speechgpt) empowers large language models with cross-modal conversational abilities by three-stage training stages, using hidden units as the discrete representation of speech. LLaSM (shu2023llasm) builds a large Chinese/English speech language model that can understand and follow instructions, through pre-training and cross-modal instruction fine-tuning stages. BLSP wang2023blsp bootstraps Language-Speech Pre-training via behavior alignment of continuation writing. SALMONN (tang2023salmonn), named from a speech audio language music open neural network, boosts large language models with generic hearing abilities with a activation tuning stage by playing with the LoRA scaling factor. Qwen-audio (chu2023qwen) scales up audio-language pre-training to cover over 30 tasks and various audio types, including human speech, natural sounds, music, and songs.
@@
Previous research on Speech Large Language Models (Speech LLMs) has primarily concentrated on executing a single speech task in response to a given instruction, while the feasibility of using a single instruction to simultaneously complete multiple and complex speech tasks has remained unexplored. The employment of multi-task instructions allows for the efficient completion of several tasks at once and improves performance by dividing complex tasks into logical, related sub-tasks, such as CoT tasks. Such capabilities also suggest the robustness and generalizability of the Speech LLM.
Our initial experiments indicate that (1) prior open-source speech LLMs underperformed in multi-task scenarios, demonstrating these models’ limited ability to generalize to complex instructions; (2) reducing the LoRA scaling factor can be beneficial for multi-task instructions, but leads to a substantial degradation of the results of training tasks (tang2023salmonn), which suggests that single and multiple tasks might benefit from distinct LoRA scaling factors; (3) there is a notable decline in performance when the model encounters unseen or diverse prompts as opposed to seen prompts (3.5% vs. 2.1%, see Section 4.3), when employing various prompts to evaluate the ASR performance of the open-source model. Consequently, we introduce a curriculum learning approach that progresses from simple to complex instruction tasks, propose a prompt-aware LoRA weight adapter which dynamically adjusts the amplitude of the LoRA output according to the instruction, and further enhance the generalization by utilizing a diverse array of prompts generated by GPT-4 across all training tasks.
@@ 3 Method
The WavLLM is optimized by maximizing the following probability:
@@
where $\bm{X}$ and $\bm{T}$ are the speech input and text prompt respectively. $\bm{Y}=[\bm{y}_{1},\bm{y}_{2},...,\bm{y}_{T_{Y}}]$ is the target text output. $\bm{\Theta}$ denotes the parameters of WavLLM. The detailed template of WavLLM’s training data can be found in Appendix D.
@@ 3.1 Model Architecture
The model architecture of our framework is shown in Figure 1, which consists of speech encoders (i.e., Whisper (radford2023robust) and WavLM (chen2022wavlm)) as well as modality adapters, a large language model (i.e., LLaMA (touvron2023llama)) and a proposed prompt adapter.
@@
In order to extract both the semantic and acoustic information in the speech, we utilize two state-of-the-art speech encoders, namely Whisper and WavLM. Whisper is trained for ASR and ST tasks in a weakly supervised fashion on a massive 680k-hour speech corpus recorded in diverse conditions, making it well suited for encoding semantic information in speech. WavLM is a predictive based self-supervised learning (SSL) pre-trained model. During its pre-training stage, WavLM mixes each utterance with signals from multiple speakers in the same batch, yet selectively predicts only the targets associated with the utterance’s original speaker. Such training method allows WavLM to better extract speaker-related acoustic information. In our work, the 32-layer transformer-based encoder of Whisper-large-v2 and WavLM-base are utilized. Both modality adapters have three components, including two 1-D convolution layers to down-sample and align the output of both encoders within the temporal domain, a down-up bottleneck adapter (houlsby2019parameter), and a final linear projector. The semantic adapter receives its input from the Whisper encoder’s output, while the acoustic adapter takes a weighted sum of the hidden states from all layers of WavLM, where the weights are learnable. The outputs of both adapters are concatenated together before feedforwarding into the linear projector.
@@
[…]
@@
where $\bm{W}_{A}\in\mathbb{R}^{1\times D}$ is the matrix of attention weights, $\bm{I}_{D}(\bm{W}_{A}\bm{o})\in\mathbb{R}^{D\times M}$ is unnormalized weights for the hidden states $\bm{o}\in\mathbb{R}^{D\times M}$. $\bm{I}_{D}\in\mathbb{R}^{D\times 1}$ and $\bm{I}_{M}\in\mathbb{R}^{M\times 1}$ are the all-ones vectors. $\bm{P}^{u}\in\mathbb{R}^{D\times K}$ and $\bm{P}^{d}\in\mathbb{R}^{K\times D}$ are up-linear projection and down-linear projection layers respectively, and GeLU is the GeLU activation function hendrycks2016gaussian. The hidden states of an attention layer equipped with adaptive LoRA are expressed by:
@@ 3.2.1 Mixed Single-Task Training Stage
In the first stage, various single-task, cross-modal, speech-text pairing datasets or text-only datasets are utilized to endow the LLM with robust capabilities in speech processing and comprehension. We freeze the parameters of LLM, WavLM, and Whisper encoder, and optimize the modality adapters, a linear layer and LoRA components.
@@
The first mixed single-task training stage involves various speech tasks, including automatic speech recognition (ASR), speech-to-text translation (ST), speaker verification (SV), emotion recognition (ER), spoken-based instruction tuning and text-based instruction tuning (IT) tasks, as well as a large mount of GPT-generated speech question answering (SQA). There are various questions within the SQA tasks, including those related to the speaker and gender, as well as continuation and summary tasks. Concurrently, these tasks draw upon multiple datasets, including LibriSpeech (Panayotov2015Librispeech) with English reading speech, AMI (carletta2005ami) with multi-talker meeting recordings, as well as Fisher (cieri2004fisher) and Switchboard (godfrey1992switchboard) corpora with 2-channel English telephony conversations. Examples of the training data and prompts used to generate data with GPT-4 can be found in the Appendix A.1 and A.3 respectively. The speech audio clips of spoken-based instruction tuning task are generated by using Microsoft Azure text-to-speech API [^2] https://azure.microsoft.com/en-us/products/ai-services/text-to-speech. The detailed task information about description, data source, and data hours can be found in Appendix F.